OpenStudy (anonymous):

If you don't brush your teeth regularly, it won't take long for large colonies of bacteria to grow in your mouth. Suppose that a single bacterium lands on one of your teeth and starts reproducing by a factor of 4 every 4 hours A) Write an equation that describes the number of bacteria b in the new colony after n hours. B) How many bacteria will be in the new colony after 7 hours, 8 hours? C) After how many hours will there be at least 1 million bacteria in the new colony?

OpenStudy (anonymous):

D) Suppose that 50 bacteria, not just 1 bacterium, land in your mouth. Write an equation that describes the number of b in this colony after n hours.

OpenStudy (anonymous):

This topic is expo/log application, growth and decay.

OpenStudy (anonymous):

our original equation is f(t) = Ae^kt to the base 0

OpenStudy (anonymous):

A to the base 0 = initial amount of t = 0 t = time k = growth constant, k>0 decay constant, k<0

OpenStudy (anonymous):

b = A*e^(kt) A=1 (one bacterium) when t=4 b=4 ln 4 = kt*ln e ln 4 = k4* ln e k= (ln4)/4

OpenStudy (anonymous):

woah, you're good!

OpenStudy (anonymous):

but can you explain to me how you got those values/equations?

OpenStudy (anonymous):

sure. what's your question?

OpenStudy (anonymous):

ok first, this might sound stupid but what is the symbol on Top of A?

OpenStudy (anonymous):

this * ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

it means multiplication.

OpenStudy (anonymous):

alright :D thanks!

OpenStudy (anonymous):

how did you get the equation?

OpenStudy (anonymous):

from you.

OpenStudy (anonymous):

"our original equation is f(t) = Ae^kt to the base 0 " not sure what you mean by 'base 0', but the rest is correct.

OpenStudy (anonymous):

cause our teacher said this is the equation: